OpenStudy (anonymous):

Indefinite integral wrt x of e^x Sin (2x)?

OpenStudy (anonymous):

1/5 * e^x(sin(2x) - 2cos(2x)) + C use euler's formula

OpenStudy (anonymous):

sin(2x) = (e^(2ix) - e^(-2ix))/2i

OpenStudy (anonymous):

is this \[e^{x}sin2x\] or is this \[e^{xsin}\]?

OpenStudy (anonymous):

you need to do parts e^xsin2x-\[e^xsin2x-\int\limits_{ }^{}e^x(2cos2x)\] = \[e^xsin 2x-e^x2cos2x-4\int\limits_{}^{}e^x(sin2x)\] so add that integral to the original one \[5\int\limits_{}^{}e^xsin2x=e^xsin2x-e^x2cos2x\] then divide by 5 and thats your final answer. kinda how you would do sec^3x

OpenStudy (anonymous):

PLUS C if your picky